# Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

### v2

@@ 4.2 Computational cost
Computational cost is the first and most critical quantity to consider since methods that are too expensive cannot be used, regardless of their quality. Figure 1 shows the average computation times for the three methods, obtained on a 16-core 3.5 GHz AMD processor equipped with 128 GB of RAM. It also shows regression lines of log cost versus $\log(N)$, marked with the regression slopes. The ARC and oracle method’s slope are both very nearly $1$ as expected. The Laplace method slope is clearly larger than $1$ and, as noted above, we curtailed the sample sizes used for that case to keep costs reasonable. Similar estimated computational costs were obtained for the other seven settings as shown in the Supplementary Material are summarized in Table 2 which shows estimated computational cost rates for all eight settings. The ARC slope is consistently very close to $1$. The oracle slope is consistently close to $1$ in imbalanced settings and somewhat bigger than $1$ in balanced ones. The Laplace slope is consistently larger than $1$. It tends to be higher for balanced settings although one of the imbalanced cases also has a large slope.
@@
Imb-Lin-Lo 0.98 0.99 1.15
Table 2: For all eight settings, this table shows the observed computational cost as a rate in $N$, fit by linear regression of the logarithm of average CPU seconds on $\log(N)$.
Figure 1: Time in seconds versus sample size N for the Imb-Nul-Hi setting for all three algorithms: Laplace, ARC and the Oracle.
@@ 4.3 Regression coefficient estimation
Next, we turn to estimation of the regression coefficients, treating the intercept differently from the others. The intercept poses a challenge because it is somewhat confounded with the random effects. For instance, if we replace $\beta_{0}$ by $\beta_{0}+\lambda$ while replacing $a_{i}$ by $a_{i}-\lambda$ then $Y_{ij}$ are unchanged. Large $\lambda$ would change $\bar{a}=(1/R)\sum_{i=1}^{R}a_{i}$ by an implausible amount that should be statistically detectable given that $a_{i}\sim\mathcal{N}(0,\sigma^{2}_{A})$. On the other hand $|\lambda|=O(\sigma_{A}/\sqrt{R})$ would be hard to detect statistically. The other regression parameters are not similarly confounded with main effects in our settings. We note that when a predictor is a categorical variable that is either a function of the row index $i$ or of the column index $j$, then it can be subject to some confounding similar to the intercept’s confounding. This was observed in Ghosh et al. (2022b).
Because of the confounding described above, we anticipate that the true MSE rate for the intercept cannot be better than $\mathcal{O}(\min\{R,C\}))$ which is $\mathcal{O}(N^{-0.53})$ in our imbalanced settings and $\mathcal{O}(N^{-0.56})$ in our balanced settings. For the other coefficients $\mathcal{O}(N^{-1})$ is not ruled out by this argument. Figures 2 and 3 reports the mean square errors for the intercept and the coefficient of the first predictor estimated for different sample sizes, for the three estimators under study in the Imb-Nul-Hi setting. We report here the plot only for the first predictor because the mean squared errors of the estimates of the seven regression coefficients were essentially equivalent. All three estimators show an MSE very close to $\mathcal{O}(N^{-1})$ for $\beta_{1}$. Where we anticipated an MSE no better than $\mathcal{O}(N^{-0.53})$ (imbalanced) or $\mathcal{O}(N^{-0.56})$ (balanced) for the intercept we saw slightly better MSEs with slopes between $-0.57$ and $-0.60$, confirming our expectation that the intercept would be harder to estimate than the regression coefficients.
Figure 2: Mean squared errors for coefficient estimates in the Imb-Nul-Hi setting. The top row is for $\beta_{0}$ and the bottom row is for $\beta_{1}$. Each method is shown with a line and solid points and fitted reference lines for the other two methods (that overlap somewhat).
@@ 4.4 Variance component estimation
Here we present the estimation errors in the variance component parameters $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. The oracle method is given the true values of these parameters and so the comparison is only between ARC and the Laplace method. For the variance parameter $\sigma^{2}_{A}$, the data only have $R$ levels $a_{1},\dots,a_{R}$. If they were observed directly, then we could estimate $\sigma^{2}_{A}$ by $(1/R)\sum_{i=1}^{R}a_{i}^{2}$ and have an MSE of $\mathcal{O}(R^{-1})$. In practice, the $a_{i}$ are obscured by the presence of the signal, the noise $\varepsilon_{ij}$ and the other random effects $b_{j}$. Accordingly, the best rate we could expect for $\sigma^{2}_{A}$ is $\mathcal{O}(R^{-1})=\mathcal{O}(N^{-\rho})$ and the best we could expect for $\sigma^{2}_{B}$ is $\mathcal{O}(C^{-1})=\mathcal{O}(N^{-\kappa})$.
@@
Figure 3 shows the MSE for estimation of $\sigma_{A}$ and $\sigma_{B}$ for ARC and Laplace methods in the Imb-Nul-Hi setting. Due to the imbalance, our anticipated rates are $\mathcal{O}(N^{-0.88})$ for $\sigma_{A}$ and $\mathcal{O}(N^{-0.53})$ for $\sigma_{B}$. The ARC method does slightly better than these rates. The Laplace method attains nearly this predicted rate for $\sigma_{B}$ but does much worse for $\sigma_{A}$. We can understand both of these discrepancies in terms of biases, described next. Figure 4 shows boxplots for the parameter estimates of $\sigma_{A}$ and $\sigma_{B}$ with the ARC and Laplace methods. We can compare the center of those boxplots to the reference line at the true parameter values and see that for the ARC estimates there is a bias decreasing at a faster rate than the width of the boxes. This explains the slightly better than predicted rates that we see for ARC. Instead the Laplace method has a substantial bias that only decreases very slowly as $N$ increases, giving Laplace a worse than expected rate, especially for $\sigma_{A}$.
In order to understand the difficulties of the Laplace approximation method we turn to the literature on Laplace approximations where the dimension of the integral grows with the sample size. The crossed random effects integral has dimension $R+C$ for $N$ observations. The largest number of parameters in any of our simulations was for the imbalanced setting with $R+C=\mathcal{O}(N^{0.88})$ and this is where the Laplace approximation had its worst performance. Section 6 of Shun and McCullagh (1995) suggests that a sufficient condition for the Laplace approximation to be reliable is, in our notation, that $R+C=o(N^{1/3})$. This is not possible in a sparse setting with $RC\gg N$. Ogden (2021) notes that Shun and McCullagh (1995) use an assumption on the growth rate of the derivatives of the exponent in the Laplacian integrand and says that assumption is not realistic for generalized linear mixed models. She then gives an analysis of the hierarchical setting which we believe is much simpler than the present crossed setting. For a hierarchical model with our row random effects but no column random effects, the likelihood is a product of $R$ one dimensional integrals. The logarithm of the $i$’th integrand is a sum of $N_{i\text{\tiny$\bullet$}}$ terms. She notes that the error in a first order Laplace approximation is $\mathcal{O}(\sum_{i}N_{i\text{\tiny$\bullet$}}^{-1})$. The smallest that could be with $N$ observations in $R$ columns is $R(N/R)^{-1}=R^{2}/N$, a quantity that is $\mathcal{O}(N^{2\rho-1})=\mathcal{O}(N^{0.76})$ in our unbalanced setting. Things improve with higher order approximations. For the definition of $k$’th order in Ogden (2021), equation (10) there gives an error bound of $\mathcal{O}(\sum_{i}N_{i\text{\tiny$\bullet$}}^{-k})$. Ogden (2021) remarks that this bound might be conservative. All of our simulations had $\sum_{i}N_{i\text{\tiny$\bullet$}}^{-1}$ increasing with $N$ but only in the imbalanced null cases did it look like the Laplace method might be inconsistent (due to bias). For the approximation of one-dimensional integrals in ARC an accurate numerical approximation was employed and we saw empirical convergence rates close to those we anticipated.
@@
Figure 3: MSEs and estimated convergence rates for $\sigma_{A}$ and $\sigma_{B}$ for ARC and Laplace methods in setting Imb-Nul-Hi. In each plot there is a reference line for the covergence rate of the other method.
@@
Figure 4: Boxplots of $\sigma_{A}$ and $\sigma_{B}$ estimates for the ARC and Laplace methods in the setting Imb-Nul-Hi. There are horizontal reference lines at the true parameter values. The ARC estimates have negative bias at small $N$ while the Laplace estimates have a strong extreme negative bias for all $N$.
@@ 4.5 Other settings
The simulation results for all eight settings, are reported in full in the Supplementary Figures S5–S44. We already discussed computational cost in eight settings based on Table 2. Here we make brief accuracy comparisons. We compare our proposed ARC method to the oracle method which is infeasible because it requires knowledge of $\sigma_{A}$ and $\sigma_{B}$ and to the Laplace method which becomes infeasible for large $N$ because it does not scale as $\mathcal{O}(N)$. In most settings and for most parameters, the oracle method was slightly more accurate than ARC. That did not always hold. For instance in setting Bal-Lin-Hi (Figure S14) we see that ARC was slightly more accurate than the oracle method for $\beta_{1}$. The figures in the Supplementary material show some cases where ARC has a slight advantage over the Laplace method and some where it has a slight disadvantage. There are a small number of cases where the Laplace method has outliers at $N=10^{3}$ that cause the linear regression slope to be questionable. These are present in setting Imb-Lin-Hi (Figure S20). However the attained MSEs at $N=10^{5}$ do not differ much between ARC and Laplace in that setting. Our conclusion is that compared to the Laplace method, ARC is scalable and robust.
The Supplementary Material Section S2 makes a comparison of ARC with the maximum pairwise likelihood estimator of Bellio and Varin (2005). The comparison is made for the Imb-Nul-Hi setting that we have been focusing on. From Section 2, the cost of pairwise composite likelihood is $\Omega(\max(N^{2}/R,N^{2}/C))=\Omega(\max(N^{2-\rho},N^{2-\kappa}))$. For the imbalanced setting this is $\Omega(N^{2-0.53})=\Omega(N^{1.47}).$ We see in Section S2 that the empirical cost of the pairwise composite likelihood grows as $N^{1.42}$, close to the predicted rate. The pairwise method attains very similar parameter estimation accuracy to ARC. The most important difference for this example is that ARC costs $\mathcal{O}(N)$ while pairwise likelihood is far more expensive and does not scale to large data sets.
@@ 5 Application to the Stitch Fix Data
In this section we illustrate our ARC method on a data set from Stitch Fix. Quoting from Ghosh et al. (2022b):
Stitch Fix is an online personal styling service. One of their business models involves sending customers a sample of clothing items. The customer may keep and purchase any of those items and return the others. They have provided us with some of their client ratings data. That data was anonymized, void of personally identifying information, and as a sample it does not reflect their total numbers of clients or items at the time they provided it. It is also from 2015. While it does not describe their current business, it is a valuable data set for illustrative purposes.
The Stitch Fix data consists of $N=5{,}000{,}000$ ratings from $R=744{,}482$ clients on $C=3{,}547$ items. The data also includes client- and item-specific covariates. In this data, the binary response $Y_{ij}$ of interest was whether customer $i$ thought that item $j$ was a top rated fit or not, with $Y_{ij}=1$ for an answer of ‘yes’. The predictor variables we used are listed in Table 3. Some of the categorical variable levels in the data had only a small number of levels. The table shows how we have aggregated them.
@@
Item fit item fit fitted
@@
Table 3: Predictors available in the Stitch Fix data. There is one block of client predictors followed by a block of item predictors.
As anticipated in Section 2, we assume missing data at random and have removed rows with any missing entries. That leaves us with $N=4{,}965{,}960$ of the original $5{,}000{,}000$ ratings. We have data from $741{,}221$ clients on $3{,}523$ items. The data are not dominated by a single row or column. The customer with the most records accounts for $N\epsilon_{R}$ records where $\epsilon_{R}\doteq 1.25\times 10^{-5}$. The item with the most records accounts for $N\epsilon_{C}$ of them with $\epsilon_{C}\doteq 2.77\times 10^{-2}$. The data are sparse because $N/(RC)\doteq 1.9\times 10^{-3}.$
In a business setting one would fit and compare a wide variety of different binary regression models in order to understand the data. Our purpose here is to study large scale probit models with including crossed random effects and so we choose just one model for illustration, possibly the first model of many that one would consider. We consider the probit model with crossed random effects whose fixed effects are specified according to the symbolic model formula:
@@
Table 4 reports 1) the maximum likelihood estimates of the regression parameters under a marginal probit model that ignores the customer and item heterogeneity, and 2) the ARC estimates for the probit model with two crossed random effects for the customers and items. The z values reported in the table are computed with the observed information for the marginal probit model and with the two-way cluster-robust sandwich estimator described in Section 3.2. The ARC estimates of the variance components are $\hat{\sigma}_{A}=0.53$ and $\hat{\sigma}_{B}=0.34$. As expected, ignoring the customer and item heterogeneity lead to a large underestimation of the uncertainty in the parameter estimates and thus in the marginal probit all the predictors are strongly significant given the very large sample size. Conversely, the cross-random effects model takes into account the sources of heterogeneity and leads us to learn that the item fit is not a significant predictor of the top rank and that items made by vegetable fibers are less likely to be ranked top compared to clothes made by artificial fibers.
While we have emphasized the naive variance estimation of the marginal probit model, Table 4 also shows the attenuation bias. The naive model systematically underestimates the absolute value of the coefficients while simultaneously overestimating their statistical significance.
Marginal probit Random-effects probit
@@
Material leather or animal --12.9 --12.99 $<0.001$-15.2 --1.57 0.116
@@
Table 4: Stitch Fix Binary Regression Results. ${}^{\text{\textdagger}}$ll the estimated predictor parameters are multiplied by 100. Nonsignificant p values are marked in bold. See Table 3 for a description of the predictors. The first level of each categorial predictor in alphabetic order is used as reference level.
@@
Figure 5 compares the two-way cluster robust sandwich standard errors with 1) the naive standard errors from a probit fit ignoring dependence between items and customers and 2) the pigeonhole nonparametric bootstrap standard errors of Owen (2007) that do not assume correct model specification, as mentioned in Section 3.2. The figure shows how closely the sandwich and pigeonhole standard errors agree, while the naive standard errors are much smaller. The naive values correspond to variances underestimated by factors ranging from 4 to just over 1240, depending on the parameter and only slightly on whether we use sandwich or pigeonhole estimates of the coefficient variances. These standard errors are reported in Table S1 of the Supplementary Material. We also computed parametric bootstrap standard errors (not shown). Those were in between the naive ones and the nonparametric ones.
Figure 5: Stitch Fix Application. Comparison between the two-way cluster robust sandwich standard errors (x axis), the naive GLM standard errors (y axis, triangles) and the Pigeonhole nonparametric bootstrap standard errors (y axis, circles). The gray reference line is the first quadrant bisector $y=x$. All standard errors are multiplied by 100.
6 Bibliographic Notes
In this section we review some related literature on the difficulties that crossed random effects bring. McCullagh (2000) points out a key difference between the crossed and nested random effects. Even in an intercept-only setting with $Y_{ij}=\mu+a_{i}+b_{j}+\varepsilon_{ij}$ and all $R\times C$ values observed, there is no correct bootstrap for the variance of the simple average of the $Y_{ij}$. Resampling rows and columns independently is mildly conservative in this case. That mild conservatism also extends to some sparse settings (Owen, 2007). Gao and Owen (2017) provide an $\mathcal{O}(N)$ algorithm to consistently estimate an intercept plus variance components. Gao and Owen (2020) provides an $\mathcal{O}(N)$ algorithm for linear regression with crossed random effects; that algorithm is less efficient than generalized least squares but the variance estimates from it properly account for all of the random effects. They also show that computing the Gaussian likelihood for crossed random effects at just one parameter value costs $\Omega(N^{3/2})$.
Ghosh et al. (2022a) provide a backfitting algorithm for generalized least squares that costs $\mathcal{O}(N)$ per iteration and they give asymptotic conditions under which the number of iterations is $\mathcal{O}(1)$. Ghandwani et al. (2023) extended the algorithm to a linear mixed model that include random slopes. Ghosh et al. (2022b) consider logistic regression with random intercepts adapting an integral approximation from Schall (1991). Their iterations cost $\mathcal{O}(N)$ and empirically the algorithm runs in $\mathcal{O}(N)$ time, but there is no proven bound on the number of iterations required.
The computational challenge from crossed random effects is also present for Bayesian methods. A plain Gibbs sampler costs $\mathcal{O}(N)$ per iteration but takes $\Omega(N^{1/2})$ iterations for the model in Gao and Owen (2017). Papaspiliopoulos et al. (2023) derived a scalable MCMC algorithm for generalized linear models with crossed-random effects that extends previous work on a collapsed Gibbs sampler described in Papaspiliopoulos et al. (2020). This Bayesian simulation method is more general than our proposal because it is not restricted to a particular type of response. Our proposal instead derives a frequentist solution to the important problem of consistent and scalable estimation of a regression model for cross-classified binary responses. Differently from the MCMC method developed by Papaspiliopoulos et al. (2023), our approach does not require the use of simulation since it is only necessary to compute a large number of one-dimensional integrals which we approximate through an adaptive quadrature carefully implemented to avoid the risk of accumulating integration errors.
Other recent work has considered variational approximations to build computational efficient fitting of generalized linear mixed models with crossed-random effects. Goplerud et al. (2023) developed a variational approximation for scalable Bayesian estimation using an appropriate relaxing of the mean-field assumption to avoid underestimation of posterior uncertainty in high-dimensions. Parallel work on frequentist inference for generalized linear models with crossed-random effects has been conducted by Xu et al. (2023) by combining variational approximations and composite likelihoods that consider row-column decomposition in a similar way to ours. This approach is particularly convenient for Poisson and gamma regression models because in this case analytical calculations allow the approximation of one-dimensional integrals that appear in the composite likelihood to be avoided. However, the consistency of the Xu et al. (2023) estimation method for binary data has not been established yet, and only a conjecture for the logistic case has been formulated by the authors. Further investigation of the inferential properties of their proposal for the case of binary data seems warranted for a thorough comparison of the two methods. Hall et al. (2020) consider message passing algorithms for generalized linear mixed models and their Section 6 includes a crossed effecs binary regression. However it has $R=10$ and $C=6$ in our notation with $3$ replicates at each $(i,j)$ pair, and it did not address scalability.

### v4

@@ 3.2 Computational cost
Figure 2 shows the average computation times for the three methods, obtained on a 16-core 3.5 GHz AMD processor equipped with 128 GB of RAM. It also shows regression lines of log cost versus $\log N$, marked with the regression slopes. The all-row-column and oracle method’s slope are both very nearly $1$ as expected. The Laplace method slope is clearly larger than $1$ and, as noted above, we curtailed the sample sizes used for that case to keep costs reasonable. If we extrapolate the Laplace cost to $N=10^{8}$, comparable to the Netflix data (Bennett & Lanning, 2007), then the cost grows past 12.9 days while the all-row-column cost grows only to about 45 minutes. The computational cost of all-row-column can be further reduced with parallel computing by distributing the approximations of the R+C one-dimensional integrals across multiple cores.
Figure 2: The left panel shows time in seconds versus sample size $N$ for the Imb-Nul-Hi setting for all three methods: Laplace, all-row-column and the Oracle. The right panel is the same plot with times displayed on the $\log_{10}$ scale.
Similar estimated computational costs were obtained for the other seven settings as summarized in Table 2 which shows estimated computational cost rates for all eight very close to $1$. The oracle slope is consistently close to $1$ in imbalanced settings and somewhat bigger than $1$ in balanced ones. The Laplace slope is consistently larger than $1$. It tends to be higher for balanced settings although one of the imbalanced cases also has a large slope.
@@
Imb-Lin-Lo 0.98 1.00 1.15
Table 2: Computational cost for all eight settings. Linear regression slopes for $\log(\text{CPU seconds})$ versus $\log(N).$
@@ 3.3 Regression coefficient estimation
Next, we turn to estimation of the regression coefficients, treating the intercept differently from the others. The intercept poses a challenge because it is somewhat confounded with the random effects. For instance, if we replace $\beta_{0}$ by $\beta_{0}+\lambda$ while replacing $a_{i}$ by $a_{i}-\lambda$ then the $Y_{ij}$ are unchanged. Large $\lambda$ would change $\bar{a}=(1/R)\sum_{i=1}^{R}a_{i}$ by an implausible amount that should be statistically detectable given that $a_{i}\sim{N}(0,\sigma^{2}_{A})$. On the other hand $|\lambda|=O(\sigma_{A}R^{-1/2})$ would be hard to detect statistically. The other regression parameters are not similarly confounded with main effects in our settings. Ghosh et al. (2022b) saw that a categorical predictor that is a function of just the row index $i$ or just the column index $j$ brings a similar confounding.
Because of the confounding described above, we anticipate that the true mean square error rate for the intercept cannot be better than $O\{\min(R,C)\}$ which is $O(N^{-0.53})$ in our imbalanced settings and $O(N^{-0.56})$ in our balanced settings. For the other coefficients $O(N^{-1})$ is not ruled out by this argument. Supplementary Material Figure S5 reports the mean square errors for the intercept and the coefficient of the first predictor estimated for different sample sizes, for the three estimators under study in the Imb-Nul-Hi setting. We report there the plot only for the first predictor because the mean square errors of the estimates of the seven regression coefficients were essentially equivalent. All three estimators show a mean square error very close to $O(N^{-1})$ for $\beta_{1}$. Where we anticipated a mean square error no better than $O(N^{-0.53})$ (imbalanced) or $O(N^{-0.56})$ (balanced) for the intercept we saw slightly better mean square errors with slopes between $-0.57$ and $-0.61$, confirming our expectation that the intercept would be harder to estimate than the regression coefficients.
@@ 3.4 Variance component estimation
Here we present the estimation errors in the variance component parameters $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. The oracle method is given the true values of these parameters and so the comparison is only between all-row-column and the Laplace method. For the variance parameter $\sigma^{2}_{A}$, the data only have $R$ levels $a_{1},\dots,a_{R}$. If they were observed directly, then we could estimate $\sigma^{2}_{A}$ by $(1/R)\sum_{i=1}^{R}a_{i}^{2}$ and have a mean square error of $O(R^{-1})$. In practice, the $a_{i}$ are obscured by the presence of the signal, the noise $\varepsilon_{ij}$ and the other random effects $b_{j}$. Accordingly, the best rate we could expect for $\sigma^{2}_{A}$ is $O(R^{-1})=O(N^{-\rho})$ and the best we could expect for $\sigma^{2}_{B}$ is $O(C^{-1})=O(N^{-\kappa})$.
@@
Figure 3 shows the mean square error for estimation of $\sigma_{A}$ and $\sigma_{B}$ for all-row-column and Laplace methods in the Imb-Nul-Hi setting. Due to the imbalance, our anticipated rates are $O(N^{-0.88})$ for $\sigma_{A}$ and $O(N^{-0.53})$ for $\sigma_{B}$. The all-row-column method does slightly better than these rates. The Laplace method attains nearly this predicted rate for $\sigma_{B}$ but does much worse for $\sigma_{A}$. We can understand both of these discrepancies in terms of biases, described next.
@@
Figure 3: Mean square errors and estimated convergence rates for $\sigma_{A}$ and $\sigma_{B}$ for all-row-column and Laplace methods in setting Imb-Nul-Hi. The dotted line in the plot for the Laplace method is the convergence rate of all-row-column. In each plot there is a reference
Figure 4 shows boxplots for the parameter estimates of $\sigma_{A}$ and $\sigma_{B}$ with the all-row-column and Laplace methods. We can compare the center of those boxplots to the reference line at the true parameter values and see that for the all-row-column estimates there is a bias decreasing at a faster rate than the width of the boxes. This explains the slightly better than predicted rates that we see for all-row-column. Instead the Laplace method has a substantial bias that only decreases very slowly as $N$ increases, giving Laplace a worse than expected rate, especially for $\sigma_{A}$.
@@
Figure 4: Boxplots of $\sigma_{A}$ and $\sigma_{B}$ estimates for the all-row-column and Laplace methods in the setting Imb-Nul-Hi. There are horizontal reference lines at the true parameter values.
@@ 3.5 Other settings
The simulation results for all eight settings, are reported in full in the Supplementary Figures S1–S24. We already discussed computational cost in eight settings based on Table 2. Here we make brief accuracy comparisons. We compare our proposed all-row-column method to the oracle method which is infeasible because it requires knowledge of $\sigma_{A}$ and $\sigma_{B}$ and to the Laplace method which becomes infeasible for large $N$ because it does not scale as $O(N)$. In most settings and for most parameters, the oracle method was slightly more accurate than all-row-column. That did not always hold. We see that all-row-column was slightly more accurate than the oracle method for $\beta_{1}$ in three of the eight settings, namely Bal-Lin-Hi (Figure S3), Bal-Lin-Lo (Figure S4) and Imb-Lin-Hi (Figure S7).
The figures in the Supplementary material show some cases where all-row-column has a slight advantage over the Laplace method and some where it has a slight disadvantage. There are a small number of cases where the Laplace method has outliers at $N=10^{3}$ that cause the linear regression slope to be questionable. These are present in the estimates of $\sigma_{A}$ and $\sigma_{B}$ for setting Imb-Lin-Hi (Figure S23). However the attained mean square errors at $N=10^{5}$ do not differ much between all-row-column and Laplace in that setting (Figure S15). Our conclusion is that compared to the Laplace method, all-row-column is scalable and robust.
The Supplementary Material §S3 makes a comparison of all-row-column with the maximum pairwise likelihood estimator of Bellio & Varin (2005). Their pairwise likelihood involves all the pairs of correlated observations, that is those pairs that share the row- or the column-random effect. The pairwise likelihood is not computationally attractive. The number of pairs of data points far exceeds the $O(N)$ constraint. In a setting with $R$ rows, the average row has $N/R$ elements in it. If all the rows had that many elements then the number of pairs would be $\Omega\{R(N/R)^{2}\}=\Omega(N^{2}/R)$. Unequal numbers of observations per row can only increase this count. As a result the cost must be $\Omega\{\max(N^{2}/R,N^{2}/C)\}=\Omega\{\max(N^{2-\rho},N^{2-\kappa})\}$, so the cost cannot be $O(N)$.
The comparison is made for the Imb-Nul-Hi setting that we have been focusing on. For that imbalanced setting, the cost of the pairwise likelihood is $\Omega(N^{2-0.53})=\Omega(N^{1.47}).$ We see in §S3 that the empirical cost of the pairwise composite likelihood grows as $N^{1.46}$, close to the predicted rate. The pairwise method attains very similar parameter estimation accuracy to all-row-column. The most important difference for this example is that all-row-column costs $O(N)$ while pairwise likelihood is far more expensive and does not scale to large data sets.
@@ 4 Application to the Stitch Fix Data
In this section we illustrate our all-row-column method on a data set from Stitch Fix. As described in Ghosh et al. (2022b):
“Stitch Fix is an online personal styling service. One of their business models involves sending customers a sample of clothing items. The customer may keep and purchase any of those items and return the others. They have provided us with some of their client ratings data. That data was anonymized, void of personally identifying information, and as a sample it does not reflect their total numbers of clients or items at the time they provided it. It is also from 2015. While it does not describe their current business, it is a valuable data set for illustrative purposes.”
The Stitch Fix data consists of $N=5{,}000{,}000$ ratings from $R=744{,}482$ clients on $C=3{,}547$ items. The data also includes client- and item-specific covariates. In this data, the binary response $Y_{ij}$ of interest was whether customer $i$ thought that item $j$ was a top rated fit or not, with $Y_{ij}=1$ for an answer of ‘yes’. The predictor variables we used are listed in Table 3. There is one block of client predictors followed by a block of item predictors. Some of the categorical variable levels in the data had only a small number of levels. The table shows how we have aggregated them.
@@
Item fit fit of the clothing item fitted
@@
Table 3: Predictors available in the Stitch Fix data.
Some of the observations with $Z_{ij}=1$ were nonetheless incomplete with a few missing entries. Deleting them left us with $N=4{,}965{,}960$ ratings from $741{,}221$ clients on $3{,}523$ items. The data are not dominated by a single row or column. The customer with the most records accounts for $N\epsilon_{R}$ records where $\epsilon_{R}\doteq 1.25\times 10^{-5}$. The item with the most records accounts for $N\epsilon_{C}$ of them with $\epsilon_{C}\doteq 2.77\times 10^{-2}$. The data are sparse because $N/(RC)\doteq 1.9\times 10^{-3}.$
In a business setting one would fit and compare a wide variety of different binary regression models in order to understand the data. Our purpose here is to study large-scale probit models with including crossed random effects and so we choose just one model for illustration, possibly the first model of many that one would consider. We consider the probit model with crossed random effects whose fixed effects are specified according to the symbolic model formula:
@@
Table 4 reports 1) the maximum likelihood estimates of the regression parameters under a naïve probit model that ignores the customer and item heterogeneity, and 2) the all-row-column estimates for the probit model with two crossed random effects for the customers and items. The random effects probit parameter estimate $\hat{\beta}$ equals the naïve probit estimate $\hat{\gamma}$ multiplied by $(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2}$. Using adaptive Gauss-Hermite quadrature we obtained estimates $\hat{\sigma}_{A}\doteq 0.53$ and $\hat{\sigma}_{B}\doteq 0.34$. Following what we learned in the simulation studies, the estimates on all rows and columns were calculated with $k=5$ quadrature nodes. An earlier conservative calculation used $k=28$ but $k=5$ gave indistinguishable results.
The z values reported in the table are computed with the observed information for the naïve probit model and with the two-way cluster-robust sandwich estimator described in §2.3. As expected, ignoring the customer and item heterogeneity leads to a large underestimation of the uncertainty in the parameter estimates and thus in the naïve probit all the predictors are strongly significant given the very large sample size. Conversely, the crossed random effects model takes into account the sources of heterogeneity and leads us to learn that the item fit is not a significant predictor of the top rank and that items made by vegetable fibers are less likely to be ranked top compared to clothes made by artificial fibers.
Naïve probit Random-effects probit
@@
Material leather or animal --12.9 --12.99 $<0.001$-15.2 --1.57 0.1160.116{0.116}0.116
@@
Table 4: Stitch Fix Binary Regression Results. ${}^{\text{\textdagger}}$ll the estimated predictor parameters are multiplied by 100. See Table 3 for a description of the predictors. The first level of each categorial predictor in alphabetic order is used as reference level.
@@
Supplementary Material Figure S29 compares the two-way cluster robust sandwich standard errors with 1) the standard errors from the naïve probit fit multiplied by $(1+\hat{\sigma}^{2}_{A}+\hat{\sigma}^{2}_{B})^{1/2}\doteq 1.18$ to report them in the fixed effects scale of the probit model with crossed random effects and 2) the pigeonhole nonparametric bootstrap standard errors of Owen (2007) that do not assume correct model specification, as mentioned in §2.3. The Figure S29 shows how closely the sandwich and pigeonhole standard errors agree. The naïve standard errors, which ignore dependence between items and customers, correspond to variances underestimated by factors ranging from 3 to 954, depending on the parameter and only slightly on whether we use sandwich or pigeonhole estimates of the coefficient variances. Thus ignoring the dependencies from correlated data makes an enormous difference here. These standard errors are reported in Table S1 of the Supplementary Material. We also computed parametric bootstrap standard errors (not shown). Those were somewhat lower than the nonparametric standard errors as expected.
In an application like that of Stitch Fix data, the typical goal is to make inference about the probability of ranking an item top for a specific customer and a specific item. Such evaluations also require estimating the customer ($a_{i}$) and item ($b_{j}$) random effects. The estimates of those random effects are a byproduct of the adaptive Gauss-Hermite quadrature used to approximate the row- and column-wise likelihoods. Figure S30 in the Supplement shows the distribution of the estimated customer and item random effects.
5 Discussion
In Theorem 2 we proved that there exists a root of the row (or column) likelihood equation that is a consistent estimator of $\tau^{2}_{A}$ (or $\tau^{2}_{B}).$ This form of consistency is commonly called Cramer consistency. It is the same notion of consistency that Jiang (2013) established for the maximum likelihood estimate. If one does not find Cramer consistency sufficient, then it is possible to construct estimators $\hat{\tau}^{2}_{A}$ and $\hat{\tau}^{2}_{B}$ that converge in probability to $\tau^{2}_{A}$ and $\tau^{2}_{B}$ as $N\to\infty$. A consistent estimator of $\tau^{2}_{A}$ can be obtained from one or more rows $i$ for which $N_{i\text{\tiny$\bullet$}}\to\infty$ and a similar approach works for $\tau^{2}_{B}$. In §S1.3 of the Supplementary Material we show how to construct such a consistent estimator and we give conditions under which the number of large rows will diverge to infinity as $N\to\infty$. See Theorem 3 there. We prefer our all-row-column estimator to an approach using just large rows, because it would be awkward to have to decide in practice which rows to use, and because we believe that there is valuable information in the other, smaller, rows.
We have used a probit model instead of a logistic one because a Gaussian latent variable is a very natural counterpart to the Gaussian random effects that are the default in random effects models. That connection simplified our modeling and computation. Gibbons & Hedeker (1997) remark that “As in the case of fixed effect models, selection of probit versus logistic response functions appears to have more to do with custom or practice within a particular discipline than differences in statistical properties.” An extension to logistic regression is outside the scope of this paper.
We are confident that our approach of combining multiple misspecified models will extend to other settings with Gaussian random effects and latent variables. The code we use already handles ordinal regression. Extensions to more than two effects or multivariate responses may well work similarly, but are outside the scope of this article.
We conclude with some additional references about recent work on inference for data with a crossed design. Goplerud et al. (2025) developed a variational approximation for scalable Bayesian estimation using an appropriate relaxing of the mean-field assumption to avoid underestimation of posterior uncertainty in high dimensions.
Xu et al. (2023) combine variational approximations and composite likelihoods that consider row-column decomposition in a similar way to ours. This approach is particularly convenient for Poisson and gamma regression models because in this case analytical calculations allow the approximation of one-dimensional integrals that appear in the composite likelihood to be avoided. In the binary case that we considered in our article, the approach of Xu et al. (2023) requires numerical integration and its consistency has not yet been established.
Hall et al. (2020) consider message passing algorithms for generalized linear mixed models and their §6 includes a crossed effects binary regression. However it has $R=10$ and $C=6$ in our notation with $3$ replicates at each $(i,j)$ pair, and it did not address scalability. Ruli et al. (2016) proposed an improved version of the Laplace approximation to overcome the potential failure of the usual Laplace approximation and also illustrates it in the case of generalized linear models with crossed random effects in their §3.5. However, this proposal is not scalable, as illustrated by the numerical results included in Ruli et al. (2016).